# On the distance between mean and geometric median in high dimensions

### v2

@@ A.1 Proofs
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ and $\mathbf{X}:=(X_{1},\ldots,X_{p})$ for $p\in\mathbb{N}$. Then, when $p\geq M+2$, there exists a unique geometric median of $\mathbf{X}$.
@@
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$. Set $\mathbf{X}:=(X_{1},\ldots,X_{p})$ for $p\in\mathbb{N}$, denote $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$ and let $\mathbf{m}_{p}$ be the geometric median of $\mathbf{X}$. Then, when $p\geq 2(M+1)+1$,
$\mathbf{m}_{p}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]$
@@
Let $G(\mathbf{m}):=\|\mathbf{X}-\mathbf{m}\|$ for $\mathbf{m}\in\mathbb{R}^{p}$. We have
$|\partial_{i}G(\mathbf{m})|=\left|\frac{(m_{i}-X_{i})}{\|\mathbf{X}-\mathbf{m}\|}\right|\leq\frac{\|\mathbf{X}-\mathbf{m}\|_{1}}{\|\mathbf{X}-\mathbf{m}\|}\leq\sqrt{p}$
by $\ell_{1}$ versus $\ell_{2}$ norm equivalence. Consequently, by [14, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
@@
Since $\mathbf{m}_{p}$ optimizes $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$, we have
$0=\operatorname{E}\left[\,\frac{\mathbf{X}-\mathbf{m}_{p}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right].$ (6)
By Lemma 2.2 and 2.3 with $k=\tfrac{1}{2}$ (both proven separately later), we have for $p\geq 2(M+1)+1$ that
$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]\leq\frac{\beta(0.5,C,\mathcal{D})}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]}}\leq\frac{\beta(0.5,C,\mathcal{D})}{\sigma_{\min}\sqrt{p}}<\infty$
by $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{2}\,\right]\geq\operatorname{E}\left[\,\|\mathbf{X}-\boldsymbol{\mu}\|^{2}\,\right]$ due to MSE optimality of the mean. Therefore, adding $0=\operatorname{E}\left[\,\tfrac{\boldsymbol{\mu}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]$ to (6), we find
$\operatorname{E}\left[\,\frac{\mathbf{m}_{p}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]=\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right].$
Pulling constants out of the expectations and dividing by $\operatorname{E}\left[\,\tfrac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]>0$ (otherwise $\mathbf{X}=\mathbf{m}_{p}$ almost surely, a contradiction to $\sigma_{\min}>0$), we have shown the claim. ∎
See 2.2
@@
By assumption, we have $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ for all $i\in\mathbb{N}$. Using the triangle inequality, optimality of the geometric median, Jensen’s inequality and the variance bound yields
$\|\mathbf{m}_{p}-\boldsymbol{\mu}\|=\|\operatorname{E}\left[\,\mathbf{m}_{p}-\boldsymbol{\mu}\,\right]\|=\|\operatorname{E}\left[\,\mathbf{m}_{p}-\mathbf{X}\,\right]\|\leq\operatorname{E}\left[\,\|\mathbf{m}_{p}-\mathbf{X}\|\,\right]\leq\\ \operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|\,\right]\leq\sqrt{\operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|^{2}\,\right]}\leq\sqrt{Cp}$
@@
See 2.3
@@
To simplify notation and stress the additive structure, we denote $S(\mathbf{v}_{p}):=\sum_{i=1}^{p}(Y_{i}-v_{p,i})^{2}=\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}$ and $s_{p}:=\operatorname{E}\left[\,S(\mathbf{v}_{p})\,\right]$ for $p\in\mathbb{N}$. The quantity of interest thus is $\operatorname{E}\left[\,(S(\mathbf{v}_{p})/s_{p})^{-k}\,\right]$.
Let $k>0$ and $p\geq\lceil 4k\rceil(M+1)+1$. We use $n:=\lfloor\tfrac{p-1}{M+1}\rfloor+1$, and note $n\geq 4k$.
We repeatedly make use of the following bounds on $s_{p}$
$\sigma_{\min}^{2}p\leq s_{p}\leq\sum_{i=1}^{p}\operatorname{E}\left[\,Y_{i}^{2}\,\right]+\operatorname{E}\left[\,\sum_{i=1}^{p}2|Y_{i}\cdot v_{p,i}|\,\right]+\|\mathbf{v}_{p}\|^{2}\leq\\ Cp+\operatorname{E}\left[\,\sum_{i=1}^{p}|Y_{i}|^{2}+(v_{p,i})^{2}\,\right]+C_{2}p\leq 2(C+C_{2})p,$ (7)
where the lower bound follows from $\operatorname{E}\left[\,(Y_{i}-v_{p,i})^{2}\,\right]\geq\operatorname{E}\left[\,(Y_{i}-\operatorname{E}\left[\,Y_{i}\,\right])^{2}\,\right]=\sigma_{\min}^{2}$, and we used Cauchy’s inequality $|ab|\leq\tfrac{a^{2}}{2}+\tfrac{b^{2}}{2}$.
For the first step of the proof, recall that the expectation of a nonnegative random variable $Y$ can be written as $\int_{0}^{\infty}\operatorname{P}\left(\,Y\geq t\,\right)\mathrm{d}t$, irrespective of whether it is finite or not. Thus
$\operatorname{E}\left[\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\,\right]=\int_{0}^{\infty}\operatorname{P}\left(\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\geq t\,\right)\mathrm{d}t=\int_{0}^{\infty}\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\mathrm{d}t.$ (8)
We split the integral in (8) into three parts, obtained by splitting the positive reals at the positions $r^{k}\leq R^{k}$, where
$r$ $:=48(C+C_{2})/\sigma_{\min}^{2}$ $R$ $:=r\cdot p^{6}.$ (9)
We bound the integrand in three different ways, depending on the segment.
Step 1: upper most segment Let $t\in(R^{k},\infty)$, meaning we need to bound the probability that $S(\mathbf{v}_{p})$ is close to zero. Since zero has a finite distance to $\operatorname{E}\left[\,S(\mathbf{v}_{p})\,\right]$, common moment-based concentration inequalities result in a constant bound as $t\to\infty$. This is prohibits their application here, since we need an integrable bound. Instead, we use the existence of densities and their uniform boundedness to obtain a bound that vanishes as $t\to\infty$.
Note that if $X$ has density $f$, then, for arbitrary $m\in\mathbb{R}$, the density of $(X-m)^{2}$ is $0.5(f(m+\sqrt{x})+f(m-\sqrt{x}))/\sqrt{x}$ with $x\in\mathbb{R}_{+}$. Consequently, for all $i\in\mathbb{N},m\in\mathbb{R}$ and $\tau\geq 0$,
$\operatorname{P}\left(\,(Y_{i}-m)^{2}\leq\tau\,\right)\leq\int_{0}^{\tau}\frac{c}{\sqrt{x}}\mathrm{d}x=2c\sqrt{\tau}.$ (10)
We aggregate this bound on an independent subset of $(Y_{i})_{i=1,\ldots,p}$. Note that by construction of $n$, it holds that $(n-1)(M+1)+1\leq p$. By $M$-dependence, $(Y_{j(M+1)+1})_{j=0,\ldots,n-1}$ are independent random variables. Thus
[…]
Indeed, this bound is integrable in $t$:
$\int_{R^{k}}^{\infty}t^{-\frac{n}{2k}}\mathrm{d}t=\frac{1}{-\frac{n}{2k}+1}\left[t^{-\frac{n}{2k}+1}\right]_{R^{k}}^{\infty}\overset{\frac{-n}{2k}+1<0}{=}\\ \frac{1}{\frac{n}{2k}-1}R^{-\frac{n}{2}+k}=\frac{1}{\frac{n}{2k}-1}p^{-3n+6k}r^{-\frac{n}{2}+k}\overset{n\geq 4k}{\leq}p^{-3n+6k}r^{-\frac{n}{2}+k}.$
Thus, the contribution of the upper most segment to the expectation of interest is
[…]
Step 2: middle segment Let $t\in(r^{k},R^{k})$. Then,
$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\leq\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right).$ (13)
The right hand side probability, i.e. that a sum of random variables deviates from its expectation, can efficiently be treated by concentration inequalities from the literature. Here, we apply a Hoeffding-type bound allowing for partial dependence among the summands. Since Hoeffding requires bounded random variables, we begin by approximating $S(\mathbf{v}_{p})$ and $s_{p}$ by quantities involving bounded random variables.
for all $1\leq i\leq p$. We set $\mathcal{I}:=\{1\leq i\leq p\mid|v_{p,i}|\leq B\}$. Note that $\mathcal{I}$ contains at least $\lfloor\tfrac{p}{2}\rfloor$ elements. Otherwise,
$\|\mathbf{v}_{p}\|^{2}\geq\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}v_{p,i}^{2}>\sum_{i\in\{1,\ldots,p\}\setminus\mathcal{I}}(\sqrt{2C_{2}})^{2}\geq\lceil\frac{p}{2}\rceil 2C_{2}\geq C_{2}p,$
a contradiction to the theorem assumptions. Defining
$A_{p,i}$ $:=(\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)-v_{p,i})^{2}$ $a_{p}$ $:=\operatorname{E}\left[\,\sum_{i\in\mathcal{I}}A_{p,i}\,\right],$
we note that by the previous discussion
$a_{p}\geq\sum_{i\in\mathcal{I}}\operatorname{Var}\left[\,\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)\,\right]\geq\lfloor\tfrac{p}{2}\rfloor\cdot\frac{\sigma_{\min}^{2}}{2}.$ (14)
Note that for $i\in\mathcal{I}$, we have $A_{p,i}\leq(Y_{i}-v_{p,i})^{2}$. Consequently, $S(\mathbf{v}_{p})\geq\sum_{i\in\mathcal{I}}(Y_{i}-v_{p,i})^{2}\geq\sum_{i\in\mathcal{I}}A_{p,i}$ and we have
$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq\frac{s_{p}}{r}\,\right)=\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-(a_{p}-\frac{s_{p}}{r})\,\right).$ (15)
We have
[…]
Recalling $r=48(C+C_{2})/\sigma_{\min}^{2}$, we have
[…]
If $p=2$, then also $\tfrac{2(C+C_{2})p}{r}\leq\lfloor\tfrac{p}{2}\rfloor\tfrac{1}{2}\tfrac{\sigma_{\min}^{2}}{2}\leq\tfrac{a_{p}}{2}$. Thus, in particular for the $p$ fixed at the beginning of the proof, by (16) it holds that
$a_{p}-\frac{s_{p}}{r}\geq\frac{a_{p}}{2}\geq\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}.$
Inserting this in (15), we obtain
$\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq\frac{s_{p}}{r}\,\right)\leq\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right).$ (17)
We now apply the Hoeffding-type bound. First, note that for $i\in\mathcal{I}$, we have $0\leq A_{p,i}\leq(2B)^{2}$ because both $|\operatorname{sign}(Y_{i})\min(|Y_{i}|,B)|\leq B$ and $|v_{p,i}|\leq B$.
The Hoeffding-type bound from [7, Thm. 2.1] builds on $\chi(\mathcal{I})$, the number of sets containing fully independent variables needed to cover $\mathcal{I}$. By $M$-dependence, we may cover $\{1,\ldots,p\}$ by $M+1$ sets of independent variables: consider the index sets $\mathcal{I}_{j}:=\{j+l(M+1):l=0,\ldots,\big\lfloor\frac{p-j}{M+1}\big\rfloor\}.$ for $j=1,\ldots,M+1$. Then $(A_{p,i})_{i\in\mathcal{I}_{j}}$ are independent for all fixed $j=1,\ldots,M+1$. Restricting this covering to $\mathcal{I}$, we obtain $\chi^{*}(\mathcal{I})\leq\chi(\mathcal{I})\leq M+1$ in the notation of [7]. Thus
$\operatorname{P}\left(\,\sum_{i\in\mathcal{I}}A_{p,i}\leq a_{p}-\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4}\,\right)\leq\exp\left(-2\frac{(\lfloor\tfrac{p}{2}\rfloor\frac{\sigma_{\min}^{2}}{4})^{2}}{(M+1)\sum_{i\in\mathcal{I}}(2B)^{4}}\right)\leq\\ \exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right).$
Inserting this into equation (17), we can bound the contribution of the middle segment to the expectation of interest as
[…]
Step 3: lower segment Let $t\in(0,r^{k})$. Since $r^{k}$ is finite and does not depend on $p$, it suffices to bound the integrated probability by $1$. Thus, the contribution of the lower segment to the expectation of interest is
$\int_{0}^{r^{k}}\operatorname{P}\left(\,S(\mathbf{v}_{p})\leq s_{p}t^{-1/k}\,\right)\mathrm{d}t\leq r^{k}.$ (19)
Final step It remains to aggregate the three segments. Inserting (12), (18), (19) into (8), we have shown
$\operatorname{E}\left[\,\left(\frac{s_{p}}{S(\mathbf{v}_{p})}\right)^{k}\,\right]\leq r^{k}\left(\frac{2c\sqrt{2(C+C_{2})}}{p\sqrt{r}}\right)^{n}p^{-3n/2+6k}+\\ r^{k}\cdot\exp\left(-\frac{\lfloor\tfrac{p}{2}\rfloor^{2}\cdot\sigma_{\min}^{4}}{8(M+1)p(2B)^{4}}\right)\cdot p^{6k}+r^{k}.$
We have shown $\sup_{p}\operatorname{E}\left[\,(S(\mathbf{v}_{p})/s_{p})^{-k}\,\right]<\infty$ if the right hand side remains bounded as $p\to\infty$. Indeed, the basis of the first term is smaller than $1$ if $p$ is large enough and the exponent is $n\geq 4k>0$; further by $n\geq 4k$ we have $3n/2+6k\leq 0$, and finally the exponential decay by $\lfloor\tfrac{p}{2}\rfloor^{2}/p$ is faster than the growth of $p^{6k}$.
Note that the right hand side of the bound directly only depends on $c,C$ and $C_{2}$, $k$, $\sigma_{\min}^{2}$ and $M$ and that indirectly $r,n,B$ are functions of $M,q,p,C,C_{2},\sigma_{\min}^{2}$ only. Thus, the finite supremum of the right hand side also only depends on those quantities. ∎
@@
Let $(Y_{i})_{i\in\mathbb{N}}$ be a sequence of real random variables such that $\sup_{i\in\mathbb{N}}\operatorname{E}\left[\,|Y_{i}|^{q}\,\right]\leq C_{1}<\infty$ for some $q>2$ and $\inf_{i\in\mathbb{N}}\operatorname{Var}\left[\,Y_{i}\,\right]\geq\sigma_{\min}^{2}$. Then there exists $\bar{B}>0$ depending only $q,C_{1}$ and $\sigma_{\min}^{2}$ such for all $B\geq\bar{B}$
@@
Let $B>0$ be arbitrary. Then, for $j\in\{1,2\}$, $i,p\in\mathbb{N}$, using Markov’s inequality, we have
@@
Further, note $\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\cdot 1_{\{|Y_{i}|\leq B\}}=Y_{i}\cdot 1_{\{|Y_{i}|\leq B\}}$. Thus, we have
@@
See 2.4

### v3

@@ A.5 Additional Lemmas
Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$. Then, when $p\geq M+2$, there exists a unique geometric median of $\mathbf{X}$.
@@
Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$. Let $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$ and let $\mathbf{m}$ be the geometric median of $\mathbf{X}$. Then, when $p\geq 2(M+1)+1$,
$\mathbf{m}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right]$
@@
Let $G(\mathbf{m}):=\|\mathbf{X}-\mathbf{m}\|$ for $\mathbf{m}\in\mathbb{R}^{p}$. We have for almost all $\mathbf{m}$ that
$|\partial_{i}G(\mathbf{m})|=\left|\frac{(m_{i}-X_{i})}{\|\mathbf{X}-\mathbf{m}\|}\right|\leq\frac{\|\mathbf{X}-\mathbf{m}\|}{\|\mathbf{X}-\mathbf{m}\|}=1.$
Consequently, by [16, Ch. 3.3, Ex. 3.3], $\operatorname{E}\left[\,G(\mathbf{m})\,\right]$ is differentiable and
@@
The geometric median $\mathbf{m}$ of $\mathbf{X}$ (see Lemma A.1) optimizes $\mathbf{m}\mapsto\operatorname{E}\left[\,G(\mathbf{m})\,\right]$, implying that
$0=\operatorname{E}\left[\,\frac{\mathbf{X}-\mathbf{m}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$ (23)
By Lemma 3.2 with $k=\tfrac{1}{2}$ (note $\|\mathbf{m}\|^{2}\leq 2(\|\boldsymbol{\mu}\|^{2}+\|\mathbf{m}-\boldsymbol{\mu}\|^{2})\leq 2(C^{2}+C)p$ by Lemma A.3; both Lemmas are proven separately later), we have for $p\geq 2(M+1)+1$ that
$\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]\leq\frac{\beta(0.5,2(C^{2}+C),\mathcal{D})}{\sqrt{\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{2}\,\right]}}\leq\frac{\beta(0.5,2(C^{2}+C),\mathcal{D})}{\sigma_{\min}\sqrt{p}}<\infty$
by $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{2}\,\right]\geq\operatorname{E}\left[\,\|\mathbf{X}-\boldsymbol{\mu}\|^{2}\,\right]$ due to MSE optimality of the mean. Therefore, adding $0=\operatorname{E}\left[\,\tfrac{\boldsymbol{\mu}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right]$ to (23), we find:
$\operatorname{E}\left[\,\frac{\mathbf{m}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right]=\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$
Pulling constants out of the expectations and dividing by $\operatorname{E}\left[\,\tfrac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]>0$ (otherwise $\mathbf{X}=\mathbf{m}$ almost surely, a contradiction to $\sigma_{\min}>0$), we have shown the claim. ∎
Lemma A.3.
Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$. Then, $\|\mathbf{m}-\boldsymbol{\mu}\|\leq\sqrt{Cp}$.
@@
By assumption, we have $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ for all $i\in\mathbb{N}$. Using the triangle inequality, optimality of the geometric median, Jensen’s inequality and the variance bound yields that
$\|\mathbf{m}-\boldsymbol{\mu}\|=\|\operatorname{E}\left[\,\mathbf{m}-\boldsymbol{\mu}\,\right]\|=\|\operatorname{E}\left[\,\mathbf{m}-\mathbf{X}\,\right]\|\leq\operatorname{E}\left[\,\|\mathbf{m}-\mathbf{X}\|\,\right]\\ \leq\operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|\,\right]\leq\sqrt{\operatorname{E}\left[\,\|\boldsymbol{\mu}-\mathbf{X}\|^{2}\,\right]}\leq\sqrt{Cp}$
@@
Lemma A.4.
Let the $p$-variate $\mathbf{X}$ belong to the class $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$. Fix $i\leq p$ and define $\check{\mathbf{X}}:=(X_{j})_{\{1\leq j\leq p\,:\,|j-i|>M\}}$ and similarly $\check{\mathbf{m}}=(m_{j})_{\{1\leq j\leq p\,:\,|j-i|>M\}}$. When $p\geq 4(M+1)$, it holds that
$m_{i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{|j-i|\leq M}(X_{j}-m_{j})^{2}}{\|\mathbf{X}-\mathbf{m}\|\|\check{\mathbf{X}}-\check{\mathbf{m}}\|(\|\mathbf{X}-\mathbf{m}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}\|)}\,\right].$
@@
Recall that by Lemma A.2 we have:
$m_{i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$ (24)
Further, $\check{\mathbf{X}}$ and $\check{\mathbf{m}}$ have at least $2(M+1)+1$ components by assumption on $p$. Using Lemma 3.3 to control $\|\check{\mathbf{m}}\|^{2}$, we may apply Lemma 3.2 with $k=\tfrac{1}{2}$ to obtain $\operatorname{E}\left[\,\tfrac{1}{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}\,\right]<\infty$. Thus:
$\operatorname{E}\left[\,\frac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}\,\right]=\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]\operatorname{E}\left[\,\frac{1}{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}\,\right]=0.$
For reals $x,y>0$ it holds that
$\frac{1}{\sqrt{x}}-\frac{1}{\sqrt{y}}=\frac{\sqrt{y}-\sqrt{x}}{\sqrt{x}\sqrt{y}}=\frac{y-x}{\sqrt{x}\sqrt{y}\,(\sqrt{x}+\sqrt{y})}.$
Subtracting $0=\operatorname{E}\left[\,\tfrac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\tfrac{X_{i}-\mu_{i}}{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}\,\right]$ from the right hand side of (24) and inserting this algebraic relation yields the claim. ∎
@@
Let $(Y_{i})_{i\in\mathbb{N}}$ be a sequence of real random variables such that $\sup_{i\in\mathbb{N}}\operatorname{E}\left[\,|Y_{i}|^{q}\,\right]\leq C_{1}<\infty$ for some $q>2$ and $\inf_{i\in\mathbb{N}}\operatorname{Var}\left[\,Y_{i}\,\right]\geq\sigma_{\min}^{2}$. Then there exists $\bar{B}>0$ depending only $q,C_{1}$ and $\sigma_{\min}^{2}$ such that
@@
Let $B>0$ be arbitrary. Then, for $j\in\{1,2\}$, $i\in\mathbb{N}$, using Markov’s inequality, we have:
@@
Further, note $\operatorname{sign}(Y_{i})\cdot\min(|Y_{i}|,B)\cdot 1_{\{|Y_{i}|\leq B\}}=Y_{i}\cdot 1_{\{|Y_{i}|\leq B\}}$. Thus, we have:
@@
Lemma A.6.
Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$. Then, for all $3\leq k\leq q$, we have that
$\operatorname{E}\left[\,|\,(X_{i}-\mu_{i})\sum_{|j-i|\leq M}(X_{j}-m_{j})^{2}\,|^{k/3}\,\right]\leq B_{4}(k,\mathcal{D})$
for all $1\leq i\leq p$ and a constant $B_{4}$ only depending on $k$ and the $\mathcal{D}$-parameters.